### v2

# TSINR: Capturing Temporal Continuity via Implicit Neural Representations for Time Series Anomaly Detection

Abstract.
Time series anomaly detection aims to identify unusual patterns in data or deviations from systems’ expected behavior. The re-construction-based methods are the mainstream in this task, which learn point-wise representation via unsupervised learning. However, the unlabeled anomaly points in training data may cause these reconstruction-based methods to learn and reconstruct anomalous data, resulting in the challenge of capturing normal patterns. In this paper, we propose a time series anomaly detection method based on implicit neural representation (INR) reconstruction, named TSINR, to address this challenge. Due to the property of spectral bias, TSINR enables prioritizing low-frequency signals and exhibiting poorer performance on high-frequency abnormal data. Specifically, we adopt INR to parameterize time series data as a continuous function and employ a transformer-based architecture to predict the INR of given data. As a result, the proposed TSINR method achieves the advantage of capturing the temporal continuity and thus is more sensitive to discontinuous anomaly data. In addition, we further design a novel form of INR continuous function to learn inter- and intra-channel information, and leverage a pre-trained large language model to amplify the intense fluctuations in anomalies. Extensive experiments demonstrate that TSINR achieves superior overall performance on both univariate and multivariate time series anomaly detection benchmarks compared to other state-of-the-art reconstruction-based methods. Our codes are available here.
time series anomaly detection, implicit neural representations, unsupervised learning
*Corresponding authors.
conference: the 31th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 3–7, 2025; Toronto, ON, Canadaccs: Computing methodologies Neural networksccs: Mathematics of computing Time series analysis
1. Introduction
Time series anomaly detection, which aims to identify unusual patterns or events across a sequence of data points collected over time (Li and Jung, 2023), has attracted a lot of attention recently. In many fields (e.g., finance, healthcare, manufacturing, and fault diagnosis), monitoring time-varying data to identify anomalies is crucial for detecting unusual behavior, potential issues, or security threats (Li and Jung, 2021; Liu et al., 2022; Chen et al., 2021; Pereira and Silveira, 2018; Li et al., 2023a). For example, in the finance field, detecting anomalous behavior in credit card data allows the prevention of theft or fraudulent transactions committed by an unauthorized party (Singh and Upadhyaya, 2012). In industrial processes, identifying anomalies helps secure safe operations, averting safety concerns and mitigating economic losses (Li et al., 2023b).
One of the major challenges for anomaly detection lies in anomalies may be rare, subtle, or have different shapes, requiring sophisticated algorithms to distinguish them from normal patterns. Moreover, time series typically exhibit trends, seasonality, and temporal dependencies, making it challenging to model such complex features. Since anomalies are typically rare and new anomalies may arise, it is difficult or expensive to collect a sufficient amount of labeled data. As one of the unsupervised-based methods, reconstruction-based methods tackle this problem by reconstructing data to learn point-wise feature representations to uncover normal patterns in the data (Zhong et al., 2024; Zhang et al., 2023). The reconstruction error, i.e., the difference between the input data and its reconstructed version, serves as a natural anomaly score. The points that have high anomaly scores are considered as anomalies, which makes it easy to interpret and understand the results.
Figure 1. (a) The diagram of INR for time series data. (b) The spectral bias property of INR to prioritize the low-frequency signals is advantageous for accomplishing time series data anomaly detection tasks.
However, it is challenging to distinguish normal and anomaly patterns because normal and anomalous points may coexist within a single instance, and anomalies may occur in the unlabeled training data (Yang et al., 2023). In addition, time series data often contains intricate patterns, and anomalies might exhibit subtle deviations from normal ones. Therefore, models may be forced to learn and reconstruct anomalous data, this makes learning a reconstruction model that can effectively capture the normal pattern challenging. This issue is also pointed out in previous work (Yang et al., 2023). Recently, implicit neural representation (INR) has become a powerful tool for continuous encoding of various signals by fitting continuous functions (Liu et al., 2023a; Molaei et al., 2023). Figure 1a depicts the diagram of INR within the context of time series data. As a continuous function, INR captures the temporal continuity of the time series, where the input is a timestamp, and the output is the corresponding value of this timestamp. In line with existing reconstruction-based anomaly detection methods, INR is also learned through a reconstruction task, making it inherently feasible for time series anomaly detection. In addition, INR possesses a spectral bias property, enabling it to prioritize the learning of low-frequency signals. Most efforts aim to alleviate this property to enhance the fitting capability for high-frequency signals (Sitzmann et al., 2020; Tancik et al., 2020). Conversely, this property is advantageous for accomplishing time series anomaly detection. As in Figure 1b, the spectral bias property enables INR to prioritize the smooth normal points and exhibit poorer performance for high-frequency abnormal data. Therefore, we aim to leverage the ability of INR to capture continuous representations and its sensitivity to anomalous data, thereby addressing the challenges of existing reconstruction-based methods.
In this paper, we propose a time series anomaly detection method based on INR reconstruction (TSINR for short). Specifically, we introduce a transformer-based architecture to predict the INR parameters of the given time series data. To better learn and reconstruct time series data, we design a novel form of continuous function to decompose time series (Cleveland et al., 1990; Fons et al., 2022). The designed function mainly comprises three components and individually learns the trend, seasonal, and residual information of time series. In addition, to further enhance the capability of INR to capture inter- and intra-channel information, we propose a group-based architecture to specifically learn the complex residual information. Simultaneously, we leverage a pre-trained large language model (LLM) to encode the original data to the feature domain, amplifying the fluctuations of anomalies from both time and channel dimensions and enabling INR to subsequently better differentiate normal and abnormal points. The major contributions of this paper are summarized as follows:
  • We utilize the spectral bias property of INR to prioritize fitting low-frequency signals and enhance sensitivity to discontinuous anomalies, thereby improving anomaly detection performance. A transformer-based architecture is employed to generate the parameters for INR, requiring only a single forward step in the inference phase.
  • We design a novel form of INR continuous function, which mainly consists of three components to implicitly learn the unique trend, seasonal, and residual information of time series. Furthermore, a group-based strategy is proposed to further learn intricate residual information.
  • We leverage a pre-trained LLM to encode the original time series to the feature domain, enabling amplification of the fluctuations of anomalies in both time and channel domains that facilitate INR to be further sensitive for noncontinuous anomaly areas. Ablation studies and visual analysis validate the aforementioned capacity to better distinguish anomaly points via our proposed framework.
  • Extensive experiments demonstrate the overall effectiveness of TSINR compared with other state-of-the-art methods on seven multivariate and one univariate time series anomaly detection benchmark datasets.
Figure 2. The overall workflow of the proposed TSINR method. The INR tokens predicted by the transformer encoder are the parameters of the INR continuous function. And the input of the INR continuous function is the timestamp $t$.
2. Related Work
2.1. Time Series Anomaly Detection
Time series anomaly detection methods primarily include statistical, classic machine learning, and deep learning methods. Statistical methods rely on analyzing the statistical properties of the data to identify patterns that deviate from the expected behavior. They are valuable for their simplicity and interpretability, but have limitations in capturing complex patterns (Cleveland et al., 1990; Coluccia et al., 2013).
Classic machine learning methods rely on manual feature extraction and various algorithms like clustering (Ruff et al., 2018; Shin et al., 2020), density estimation (Breunig et al., 2000; Yairi et al., 2017; Zong et al., 2018), and isolation forests (Liu et al., 2008) to identify anomalies in structured data. However, because they require manual feature extraction and selection, they can be labor-intensive and less effective at capturing complex patterns in data.
Deep learning methods automatically learn the features of data through deep neural networks without the need for manual intervention, and are adept at handling high-dimensional, unstructured data. They can be broadly categorized into supervised and unsupervised learning algorithms. Supervised methods are trained with labels to learn and classify both normal and anomalous behavior in the given time series data, such as NFAD (Ryzhikov et al., 2021) and MultiHMM (Li et al., 2017). However, annotating data is challenging due to the rarity of anomalies and the emergence of new anomalies. This makes it difficult to achieve effective labeling, leading to limitations in the performance of supervised methods in detecting anomalies.
On the contrary, unsupervised methods distinguish anomalous points from normal ones without relying on prior knowledge. The unsupervised methods mainly comprise 2 categories: forecasting-based methods and re-construction-based methods. Forecasting-based methods train a model to predict future values based on past observations and then identify anomalies by comparing the actual values with the predicted ones, such as ARIMA (Yaacob et al., 2010) and Telemanom (Hundman et al., 2018). Reconstruction-based methods learn and reconstruct the input data and then identify anomalies based on the difference between the original and reconstructed data, such as LSTM-VAE (Park et al., 2018), BeatGAN (Zhou et al., 2019), OmniAnomaly (Su et al., 2019), and TranAD (Tuli et al., 2022).
Recently, some methods have attempted to build a universal framework for various time series data tasks, such as FPT (Zhou et al., 2024) and TimesNet (Wu et al., 2022). Among these works, models designed for prediction tasks also perform well in anomaly detection tasks, such as DLinear (Zeng et al., 2023), ETSformer (Woo et al., 2022) and LightTS (Zhang et al., 2022). These methods demonstrate outstanding performance in learning time series features, thereby it makes sense to use these methods as baselines.
2.2. Implicit Neural Representations
Presently, INR (Xie et al., 2022) stands as a scorching topic in the domain of deep learning. It aims to learn a continuous function, often embodied as a neural network, for data representation. In this function, the input comprises coordinates, while the output consists of corresponding data values. INR learns continuous representations and has been widely applied in numerous scenarios, such as 2D image generation (Qin et al., 2022; Strümpler et al., 2022; Zhou et al., 2021a) and 3D scene reconstruction (Ran et al., 2023; Kohli et al., 2020; Guo et al., 2022), physics-informed problems (Raissi et al., 2019; Pfrommer et al., 2021) and video representation (Chen et al., 2022; Lu et al., 2023; Mai and Liu, 2022).
Current methods for learning INR parameters are primarily based on either meta-learning (Liu et al., 2023a) or feed-forward networks (Chen and Wang, 2022; Zhang et al., 2024). The key difference is that meta-learning requires a few training steps for each unseen test data, while feed-forward networks directly make predictions in a single forward pass. In this paper, we use a transformer-based architecture to generate INR parameters and it requires only a single forward in the inference phase (Chen and Wang, 2022).
In addition, there exists a phenomenon known as spectral bias (Rahaman et al., 2019), where INR tends to prefer fitting the low-frequency components of the signal (Liu et al., 2023a). Since this characteristic can affect the ability of INR to model high-frequency data, most efforts are directed towards mitigating this effect (Sitzmann et al., 2020; Tancik et al., 2020; Liu et al., 2023b). In contrast, in time series anomaly detection, this property turns out to be advantageous. In time series data, normal points exhibit relative smoothness, whereas anomalous points possess strong discontinuity. Hence, we leverage this property of INR to prioritize fitting normal data with low frequencies, making it more sensitive to anomalous data.
2.3. Implicit Neural Representations on Time Series Data
Currently, there are some studies discussing the possibility of utilizing INR for time series representation. HyperTime (Fons et al., 2022) leverages INR to learn a compressed latent representation for time series imputation and generation. TimeFlow (Naour et al., 2023) uses INR to capture continuous information for time series imputation and forecasting. In addition, the potential of employing INR for time series anomaly detection has not been fully explored yet. Only INRAD (Jeong and Shin, 2022) attempts to adopt INR to represent and reconstruct time series data to identify anomalies. INRAD aims to utilize INR to overcome deep learning limitations, like complex computations and excessive hyperparameters (e.g., sliding windows). However, it requires training an INR network for each unseen time series data in test set, leading to additional training time and inefficiency in practical applications.
Different from INRAD, our method uses the spectral bias property of INR to mitigate the impact of unlabeled anomalies for the reconstruction model. With the Transformer integration, it can detect anomalies in unseen test data without retraining, enhancing efficiency for practical applications. In addition, compared to INRAD and other INR methods, our approach incorporates several specialized designs for time series anomaly detection. Firstly, we devise a novel form of INR continuous function to capture trend, seasonal, and residual information to address unique temporal patterns. Secondly, to handle multivariate time series, we introduce a group-based architecture to bolster the representational capacity of INR. Lastly, we leverage LLM to enhance anomaly detection by amplifying anomaly fluctuations, thereby boosting the sensitivity of INR to anomalies.
3. Methodology
In this paper, we propose TSINR, a novel time series anomaly detection method based on INR reconstruction. The core idea is to leverage the spectral bias phenomenon of INR to prioritize fitting smooth normal points, thereby enhancing sensitivity to discontinuous anomalous points. In this section, we present the problem statement and introduce the overall architecture, followed by the form of INR continuous function designed for time series data and the frozen pre-trained LLM encoder applied to amplify the fluctuations of anomalies from both time and channel dimensions. The anomaly criterion is demonstrated finally.
3.1. Problem Statement
Consider a time series $X$ with $T$ timestamps: $X=\left(x_{1},x_{2},\cdots,x_{T}\right)$, where $x_{t}\in\mathbb{R}^{d}$ is the data point observed at a certain timestamp $t$ ($t\in\{1,2,\ldots,T\}$) and $d$ denotes the number of the data variables (i.e., data dimensionality). For a multivariate data, $d>1$. And for an univariate case, $d=1$. Given unlabeled input time series data $X_{train}$, for any unknown time series data $X_{test}$ with the same data dimensionality $d$ as $X_{train}$, we aim to predict $y_{test}=\left(y_{1},y_{2},\cdots,y_{T^{\prime}}\right)$, where $T^{\prime}$ is the length of $X_{test}$. And $y_{t^{\prime}}\in\{0,1\}$ denotes whether the data point is normal ($y_{t^{\prime}}=0$) or abnormal ($y_{t^{\prime}}=1$) at the certain timestamp $t^{\prime}$ ($t^{\prime}\in\{1,2,\ldots,T^{\prime}\}$).
3.2. Overall Workflow
Figure 2 shows the overall workflow of the proposed TSINR method. We employ a feed-forward transformer-based architecture to directly predict the whole weights of the INR of the given time series data (Chen and Wang, 2022). Unlike meta-learning based on gradient descent (Liu et al., 2023a), our method requires only a single forward step in the inference phase. Following a strategy similar to other transformer-based methods (Dosovitskiy et al., 2020; Zeng et al., 2023), the input time series data is normalized and segmented into patches. A frozen pre-trained LLM encoder is applied to map the input data into the feature domain to amplify the fluctuations of anomalies. Then the obtained features are converted to data tokens using a fully connected (FC) layer. Simultaneously, we initialize the corresponding INR tokens, which are learnable vector parameters. These data tokens and initialized INR tokens are fed together into a transformer encoder, which mainly consists of self-attention modules and feed-forward modules. In this transformer encoder, the knowledge interacts with data tokens and INR tokens. The learned INR tokens are mapped to the INR weights through FCs, denoted as FC. These INR weights form our INR continuous function, which is specifically designed for time series data. The designed function takes a batch of timestamps $\{t\}^{T}_{i=1}$ as input, implicitly learns the trends, seasonality, and residual information of the given time series data, and finally reconstructs the input signal. The details of the designed form of INR continuous function and the applied frozen pre-trained LLM encoder module are demonstrated in Section 3.3 and Section 3.4. And the anomaly criterion is introduced in Section 3.5. Relevent pseudo-codes are provided in Appendix A.
3.3. Form of INR Continuous Function
As shown in Figure 2, we innovatively propose a INR continuous function to better learn and reconstruct time series data. Inspired by classical time series decomposition methods (Cleveland et al., 1990; Fons et al., 2022), the proposed INR continuous function $f$ consists of three components, including trend $f_{tr}$, seasonal $f_{s}$, and residual $f_{r}$:
(1) $f(t)=f_{tr}(t)+f_{s}(t)+f_{r}(t).$
Figure 3. A sample of the proposed group-based architecture. It contains 2 global layers and 2 group layers. In this case, each group corresponds to one variable.
The trend component captures the underlying long-term patterns and focuses on slowly varying behaviors. In order to model this monotonic function, a polynomial predictor is applied (Oreshkin et al., 2019; Fons et al., 2022):
(2) $f_{tr}(t)=\sum_{i=0}^{p}\boldsymbol{w}_{tr}^{(i)}t^{i},$
where $\boldsymbol{w}_{tr}^{(i)}$ is the polynomial coefficients corresponding to the $i^{th}$ degree and $\boldsymbol{w}_{tr}^{(i)}$ is predicted by a FC network. In addition, $p$ denotes the polynomial degree and is set to be small in order to model the low-frequency information and mimic the trend.
The seasonal component grasps the regular, cyclical, and recurring short-term fluctuations. Therefore, a periodic function is employed based on Fourier series (Oreshkin et al., 2019; Fons et al., 2022):
(3) $f_{s}(t)=\sum_{i=0}^{\lfloor T/2-1\rfloor}\left(\boldsymbol{w}^{(i)}_{s}\cos(2\pi it)+\boldsymbol{w}^{(i+\lfloor T/2\rfloor)}_{s}\sin(2\pi it)\right),$
where $\boldsymbol{w}_{s}$ are Fourier coefficients learned by a FC network. This component is then able to model the periodic information and simulate typical seasonal patterns.
The residual component aims to represent the unexplained variability in the data after accounting for the trend and seasonal components. In order to capture this complex and non-periodic information, we design a group-based architecture as shown in Figure 3. For any given timestamp $t$, we design $M$ global layers and $N$ group layers with $k$ groups. The global layers capture the inter-channel information while the group layers focus on intra-channel information. The equation for calculation within global layers can be defined as:
(4) $q_{m+1}=\mathrm{ReLU}\left(\boldsymbol{w}^{(m)}_{r}q_{m}+\boldsymbol{b}^{(m)}_{r}\right),$
where $m\in[0,M)$, and $q_{m}$, $\boldsymbol{w}^{(m)}_{r}$, $\boldsymbol{b}^{(m)}_{r}$ denote the outputs, weights, and biases of the $m^{th}$ global layer respectively. For the group layers, we clone the parameters of $q_{m}$ for $n$ copies and get $\{q_{M,i}\}_{i=1}^{n}$, each of which is served as the input of the group layers of a group. Then the equation of the output of the $i^{th}$ group at the $(l+1)^{th}$ group layer is given as:
(5) $q_{M+l+1,i}=\mathrm{ReLU}\left(\boldsymbol{w}^{(M+l,i)}_{r}q_{M+l,i}+\boldsymbol{b}^{(M+l,i)}_{r}\right),$
where $l\in[0,N)$, and $q_{M+l}$, $\boldsymbol{w}^{(M+l)}_{r}$, $\boldsymbol{b}^{(M+l)}_{r}$ denote the outputs, weights, and biases of the $l^{th}$ group layer respectively. Finally, the outputs of the $N^{th}$ group layers are concatenated:
(6) $f_{r}(t)=q_{M+N,1}\oplus q_{M+N,2}\oplus\ldots\oplus q_{M+N,n}.$
3.4. Frozen Pre-trained LLM Encoder
To further enhance the ability of TSINR to detect anomalies, we leverage the representational capability of LLM. A pre-trained LLM is employed as the encoder, which has been demonstrated to process time-series data and provide cross-modal knowledge (Zhou et al., 2024). With this pre-trained LLM, we map the input data into the feature domain to amplify the fluctuations of anomalies from both time and channel dimensions. On the one hand, in the time dimension, we observe that the extracted feature of LLM involves more intense fluctuations during the anomaly interval. On the other hand, in the channel dimension, other channels have the same anomaly interval due to the ability of LLM to extract and fuse the inter-channel information. Therefore, TSINR can exhibit greater sensitivity to anomalous data, thereby enhancing its ability for anomaly detection. The corresponding experimental results and analysis can be found in Section 4.4.2.
More specifically, the self-attention layers and the feed-forward layers are frozen to preserve the prior knowledge in the pre-trained model. For any given time series data $X\in\mathbb{R}^{d\times T}$, the pre-trained LLM encoder maps it to feature domain:
(7) $Z=Encoder(X),$
where $Z\in\mathbb{R}^{d\times T}$ denotes the feature corresponding to $X$.
3.5. Anomaly Criterion
Following previous reconstruction-based anomaly detection approaches (Sakurada and Yairi, 2014; Park et al., 2018; Zhou et al., 2024), we use the reconstruction error as the anomaly score for each time series data point. The anomaly score at timestamp $t$ is defined as follows:
(8) $AnomalyScore(t)=\frac{1}{d}\sum_{i=1}^{d}\left\|x_{t,i}-f(t)_{i}\right\|^{2}.$
Based on this point-wise anomaly score, we use a parameter threshold $\delta$ to determine whether the point is abnormal or normal:
(9) $y_{t}=\begin{cases}1:\text{ abnormal }&\text{AnomalyScore}\left(t\right)\geq\delta\\ 0:\text{ normal }&\text{AnomalyScore}\left(t\right)<\delta\end{cases}.$
The threshold $\delta$ is set to label a proportion $\gamma$ of the test dataset as anomalies. And $\gamma$ is a hyper-parameter based on actual datasets.
4. Experiments
4.1. Datasets
We use eight anomaly detection benchmarks from real-world scenarios to validate the performance of our proposed method, including seven multivariate datasets (SMD (Su et al., 2019), PSM (Abdulaal et al., 2021), SWaT (Mathur and Tippenhauer, 2016), MSL (Hundman et al., 2018), SMAP (Hundman et al., 2018), PTB-XL (Jiang et al., 2023; Wagner et al., 2020), and SKAB (Katser and Kozitsin, 2020)) and one univariate dataset (UCR (Wu and Keogh, 2021)). More details are in Appendix B.
4.2. Baselines and Experimental Settings
We compare our proposed method with 11 state-of-the-art deep learning approaches, including both general frameworks designed for time series modeling and algorithms specifically tailored for time series anomaly detection: FPT (Zhou et al., 2024), TimesNet (Wu et al., 2022), ETSformer (Woo et al., 2022), FEDformer (Zhou et al., 2022), LightTS (Zhang et al., 2022), DLinear (Zeng et al., 2023), Autoformer (Wu et al., 2021), Pyraformer (Liu et al., 2021), AnomalyTransformer (Xu et al., 2021), Informer (Zhou et al., 2021b) and Transformer (Vaswani et al., 2017). Also, we compare our TSINR method with other classical anomaly detection methods and the comparison results are in Appendix F. Additionally, we employ the commonly used metrics of precision, recall, and F1-score for evaluation.
The implementation details and the default hyper-parameters are summarized here. For a fair comparison, we only employ the classical reconstruction error across all baseline models. Also, we adopt identical data processing methods and the corresponding parameter configurations. We employ the sliding window approach and use a fixed window size of 100 for all datasets. The proportion $\gamma$ mentioned in Section 3.5 is set to 0.5 for SMD dataset, 0.1 for UCR dataset, 10 for SKAB dataset, and 1 for others. These parameters adhere to the settings of previous work (Zhou et al., 2024; Xu et al., 2021). Ablation studies on the anomaly proportion $\gamma$ are in Appendix E. For the main results, our TSINR model involves 3 global layers and 2 group layers in the residual block. And the hidden dimensions are 64 and 32 respectively. The transformer encoder has 6 blocks. We use GPT2 (Radford et al., 2019) as the pre-trained LLM encoder and 6 blocks are utilized following the same settings as in FPT (Zhou et al., 2024). The experiments are conducted using the ADAM optimizer (Kingma and Ba, 2014) with an initial learning rate of $10^{-4}$. A single NVIDIA Tesla-V100 32GB GPU is applied for each dataset. And the efficiency analysis is in Appendix H.
Table 1. The overall results on multivariate and univariate datasets. The precision (P), recall (R), and F1-score (F1) values are reported, all in percentage (%). The best results of F1 on each dataset and the average P, R F1 among all datasets are in Bold and the second best ones are underlined.
Dataset Metric Trans. FED. Anomaly. Auto. Pyra. In. ETS. LightTS Dlinear TimesNet FPT TSINR
SMD P 78.32 78.45 78.72 78.49 78.49 78.37 86.63 87.04 87.27 87.98 87.27 83.09
R 65.24 65.08 65.43 65.13 65.53 65.23 75.35 78.39 80.99 81.54 81.08 80.46
F1 71.19 71.14 71.46 71.19 71.43 71.20 80.68 82.49 84.01 84.64 84.06 81.76
PSM P 90.75 99.99 98.76 99.99 99.62 99.68 98.17 98.29 98.66 98.51 98.55 99.21
R 54.68 81.89 83.25 78.99 88.46 83.30 91.36 93.60 94.70 96.27 95.79 89.37
F1 68.24 90.04 90.35 88.26 93.71 90.75 94.64 95.89 96.64 97.38 97.15 94.04
SWaT P 99.67 99.95 99.73 99.96 99.71 99.64 92.01 92.36 92.25 92.14 92.12 99.31
R 68.93 65.56 68.07 65.56 68.05 68.96 93.33 93.32 93.10 93.09 93.06 72.32
F1 81.50 79.19 80.91 79.19 80.90 81.51 92.67 92.84 92.68 92.61 92.59 83.69
MSL P 90.58 90.69 89.78 90.66 90.64 90.63 86.89 89.17 89.68 89.55 82.03 83.57
R 74.65 75.48 73.66 75.22 74.76 74.96 67.78 73.64 75.31 75.29 82.01 85.40
F1 81.85 82.39 80.93 82.22 81.94 82.06 76.16 80.66 81.87 81.80 82.02 84.47
SMAP P 90.87 89.98 90.14 90.72 89.51 90.66 90.75 90.02 89.89 89.92 90.91 91.67
R 61.44 55.89 54.00 62.58 54.59 61.69 54.68 53.90 54.01 56.56 61.01 76.42
F1 73.31 68.95 67.54 74.07 67.82 73.43 68.24 67.43 67.48 69.44 73.02 83.35
PTB-XL P 56.89 48.36 56.00 49.12 50.22 57.41 62.84 66.38 62.95 67.60 71.85 58.35
R 29.99 27.60 31.50 27.53 23.85 25.43 28.45 16.46 13.95 14.47 24.52 35.00
F1 39.28 35.14 40.32 35.28 32.34 35.25 39.17 26.38 22.84 23.84 36.57 43.75
SKAB P 87.56 86.88 91.83 87.51 89.55 88.67 85.38 83.83 86.01 85.65 86.18 89.98
R 86.72 77.71 95.04 91.10 97.27 97.27 100.00 82.01 100.00 100.00 99.21 98.65
F1 87.14 82.04 93.41 89.27 93.25 92.77 92.12 82.91 92.48 92.27 92.24 94.11
UCR P 41.13 32.96 44.79 42.82 42.12 43.97 40.12 37.70 34.55 33.11 41.00 67.29
R 33.61 25.73 34.83 33.97 35.13 35.16 29.85 29.01 29.06 29.18 32.51 62.35
F1 34.50 27.09 36.51 35.52 36.02 36.41 31.94 30.82 29.67 29.81 34.33 62.46
Average P 79.47 78.41 82.22 79.91 79.98 81.13 80.35 80.60 80.16 80.56 81.24 84.06
R 59.41 59.37 63.22 62.51 63.46 64.00 67.60 65.04 67.64 68.30 71.15 75.00
F1 67.13 67.00 70.18 69.38 69.68 70.42 71.95 69.93 70.96 71.47 74.00 78.45
4.3. Main Resutls
We compare our method with 11 other state-of-the-art approaches and the results are shown in Table 1. These results show that our method achieves superior overall performance on these benchmark datasets. These experimental results confirm that TSINR, in both multivariate and univariate scenarios, effectively captures temporal continuity and precisely identifies discontinuous anomalies. The findings affirm the robustness of TSINR across diverse datasets and showcase its potential for broader applications in diverse domains.
In multivariate scenarios, we observe that despite both MSL and SMAP being collected from NASA Space Sensors, TSINR achieved significantly greater improvements on the SMAP dataset compared to other methods. This could be attributed to the presence of more point anomalies in the SMAP dataset. Point anomalies exhibit poorer continuity compared to other anomaly patterns due to their isolated nature, representing single data points that significantly deviate from the surrounding pattern. This aligns with the property of spectral bias, making our model more sensitive to point anomalies, thereby achieving greater improvements on the SMAP dataset. Meanwhile, our approach shows moderate performance on the SWAT dataset. This is due to the poorer continuity of the SWaT dataset, which affects the fitting ability of TSINR. More detailed analysis can be found in Appendix C.
Table 2. Ablation studies on the decomposition components and the group-based architecture. The F1 score is reported and the best results are in Bold.
Decomposition Group-based SMD PSM SWaT MSL SMAP PTB-XL SKAB
78.52 92.61 81.89 82.02 73.31 40.11 93.13
80.10 93.08 82.28 82.28 78.94 40.76 94.01
79.24 93.04 82.16 82.95 78.66 42.65 93.96
81.76 94.04 83.69 84.74 83.35 43.75 94.11
Table 3. Ablation studies on the pre-trained LLM encoder. The F1 score is reported and the best results are in Bold.
Pre-trained LLM SMD PSM SWaT MSL SMAP PTB-XL SKAB UCR
80.29 92.69 82.33 83.27 79.35 40.04 93.91 62.46
81.76 94.04 83.69 84.47 83.35 43.75 94.11 60.41
Figure 4. The visualization of the original data and the corresponding features obtained from the frozen pre-trained LLM encoder. The intense fluctuations in the anomaly area are amplified from both time and channel domains.
The situation is more complex in the UCR dataset. The UCR dataset comprises 250 univariate sub-datasets, and we report the average scores. The results demonstrate that our approach still outperforms other methods in overall performance by a significant margin. These sub-datasets originate from various domains, demonstrating the generalization capability of our method. In addition, it proves the strong ability of TSINR for modeling and identifying anomalies in univariate scenarios and underscores the importance of the spectral bias constraint for anomaly detection. We compile the performance of models in each domain of the UCR dataset, with the complete experimental results presented in Appendix G.
4.4. Ablation Studies
Figure 5. The visualization of the UCR original data and the corresponding features obtained from the frozen pre-trained LLM encoder of (a) anomaly area and (b) normal area.
4.4.1. Analysis of the Decomposition Components and the Group-based Architecture
In this section, we analyze the effectiveness of the proposed decomposition components and group-based architecture. The decomposition components indicate the three components (i.e., trend, seasonal, and residual) designed in our paper. And the group-based architecture is proposed for the residual block.
The main purpose of the decomposition components is to extract the unique trend and seasonal information of the time series data. The results in Table 2 indicate that capturing these distinctive features significantly enhances the capability for anomaly detection. In addition, the group-based architecture is designed to enhance the representational capacity of INR for multivariate data. Experimental results indicate an improvement in the capability for anomaly detection when employing the proposed group-based architecture. This is because modeling multiple variables and capturing both inter- and intra-channel information with a simple continuous function, which only consists of fully-connected layers, is challenging. Our approach addresses this by dividing the variables into several groups and applying independently fully-connected layers in different groups, thereby reducing the number of variables each group needs to model and improving the representational capacity. The global layers extract the inter-channel information, while the group layers selectively focus on detailed information for specific channels. This enhances the representational capability for each variable without losing any knowledge. Detailed ablation studies on the number of groups are left in Appendix D.
4.4.2. Analysis of Pre-trained LLM Encoder
Figure 6. The visualization of the original data, the TSINR reconstructed data, and the corresponding anomaly score. In the anomaly area, anomaly scores significantly increase, demonstrating the sensitivity of TSINR to anomalies.
Figure 7. The visualization of the ground-truth anomalies and anomaly scores of TSINR for different types of anomalies.
Further, we prove the validity of the pre-trained LLM encoder, which is utilized to encode the data into the feature domain to amplify the fluctuations of anomalies and thereby enhance the capability of TSINR in identifying anomalies. Table 3 displays the ablation studies of the pre-trained LLM encoder. For the multivariate datasets, it can be observed that applying this encoder enhances the performance of anomaly detection. To further demonstrate the effectiveness, we compare the raw data with the features extracted through the encoder. As shown in Figure 4, the figures in the first row illustrate that during the time interval when anomalies occur, the extracted features exhibit more pronounced fluctuations compared to the original data. This implies that the discontinuity in anomalies is increased in time domain. Also, these extracted features incorporate inter-channel information, providing a manifestation of anomalies among all variables. As shown in the second line, the features exhibit anomalous fluctuations in the same time interval as other channels, whereas the original data only shows a brief peak. This verifies that the anomalies are shared in channel domain. Based on these results, we indicate that utilizing the pre-trained LLM encoder can effectively enhance abnormal information both intra- and inter-channel. This aligns with the spectral bias of INR, making our model more sensitive to anomalous data.
In contrast, on the UCR dataset, using the pre-trained LLM encoder actually decreased the performance of anomaly detection. This is because the advantages observed in the aforementioned multivariate datasets do not apply to the UCR dataset. On the one hand, there are no anomalies in the UCR training data. Consequently, during inference, the LLM encoder still extracts features according to the normal pattern. As shown in Figure 5, the highly discontinuous anomalous data becomes relatively smooth after feature extraction. Meanwhile, there is little variation in features between normal and abnormal areas. This prompts TSINR to fit the anomalous data, thereby reducing its sensitivity to anomalies. On the other hand, the UCR dataset involves only a single variable, thus the inter-channel information provided by the LLM encoder is meaningless.
4.5. Visual Analysis
In order to demonstrate that our approach is sensitive to discontinuous abnormal data, we compare the original data with the reconstructed values of the TSINR model. As shown in Figure 6, the smooth normal points are well-fitted, while the discontinuous abnormal points are not. The anomaly scores significantly increase when anomalies occur, which aids the TSINR model in distinguishing abnormal points, demonstrating its sensitivity to the anomaly points. It is worth noting that, despite the small and highly fluctuating values of anomaly scores on the UCR dataset, this still holds meaningful significance. The UCR dataset involves only one anomaly, hence even minor fluctuations are helpful for the model to pinpoint these anomaly points.
Furthermore, we validate the robustness of TSINR with the synthetic data generated for time series anomaly detection. It has univariate time series and involves different types of anomalies (Yang et al., 2023; Lai et al., 2021), including the point-wise anomaly (global point and contextual point anomalies) and pattern-wise anomalies (shapelet, seasonal, and trend anomalies). As shown in Figure 7, the red points are anomaly points and the red areas are anomaly areas. It can be seen that TSINR can robustly detect various types of anomalies from normal points with relatively high anomaly scores.
5. Conclusion
Time series data anomaly detection plays a pivotal role in ensuring the reliability, security, and efficiency of systems across various domains. Reconstruction-based methods are mainstream approaches for this task because they do not require label information and are easy to interpret the detection results. However, unlabeled anomalous points in the training data can negatively impact the performance of these reconstruction models. To address this issue, this paper proposes a novel algorithm named TSINR for time series anomaly detection based on INR reconstruction. We utilize the spectral bias of INR to prioritize fitting continuous normal data and capture temporal continuity, thus enhancing sensitivity to discontinuous anomalous data. A transformer-based architecture is employed to predict the parameters of INR. To cope with the complex patterns of time series data, we specifically design a formulation of continuous function. It aims to implicitly learn the trend, seasonal, and residual information and capture the inter- and intra-channel information of the time series data. Besides, we leverage a frozen pre-trained LLM encoder to map the original data to the feature domain, thus amplifying the fluctuations of anomalies from both time and channel domains and enabling TSINR to better distinguish between abnormal and normal points. Experimental results indicate that TSINR exhibits superior overall performance on both multivariate and univariate benchmark datasets compared to other state-of-the-art algorithms. Also, ablation studies verify the effectiveness of each component, and visual analysis demonstrates the sensitivity of TSINR to anomalous points. In this work, we demonstrate the potential of INR in time series data tasks. In future work, we plan to explore the performance of INR on other time series tasks, including imputation, classification, and long-term and short-term forecasting. We believe that INR has the potential to become a unified framework for various time-series data tasks.
Appendix
Appendix A Algorithm
Pseudo-codes of the main components of TSINR are presented in 1. All experiments are implemented based on PyTorch. The proposed form of INR continuous function takes the timestamp $t$ as input, which individually passes through three blocks: trend, seasonal, and residual, to predict the corresponding time series information. Then, the predictions from these blocks are summed up to obtain the reconstructed time series data as shown in Eq. 1. For the trend and seasonal components, we use two single-layer MLPs to fit the trend information in Eq. 2 and the periodic information in Eq. 3 separately. For the residual component, we propose a group-based architecture to learn intricate residual information. A 5-layer MLP is utilized for learning complex residual information in Eq. 4 and Eq. 5, comprising 3 global layers and 2 group layers. Finally, we use a transformer-based architecture to generate the INR parameters. The original data is encoded by a pre-trained LLM encoder and the obtained features are passed into the transformer encoder as data tokens. The INR tokens are learned through transformer encoders and further transformed into parameters for the MLPs in the trend, seasonal, and residual blocks through the post functions.
Algorithm 1 TSINR Anomaly Detection Algorithm
1:Input: A batch of time serial data $X\in\mathbb{R}^{B\times d\times T}$, trained initialized INR token $I_{init}\in\mathbb{R}^{B\times N_{I}\times C}$, the timestamp $C\in R^{B\times T}$, the anomaly threshold $\delta$
2:Output: The predicted anomaly label $Y\in[0,1]^{B\times T}$
3:Normalize and patch $X$ to data feature $D\in\mathbb{R}^{B\times N_{d}\times C}$
4:Concatenate $D$ and $I_{init}$ to feature $F\in\mathbb{R}^{B\times(N_{d}+N_{I})\times C}$
5:Encode $F$ with LLM encoder and learnable Transformer encoder to get processed feature $F^{\prime}\in\mathbb{R}^{B\times(N_{d}+N_{I})\times C}$
6:Obtain the data-dependant INR token $I^{\prime}\in\mathbb{R}^{B\times N_{I}\times C}$ by choosing the corresponding part of $F^{\prime}$
7:Using FC layer to convert I’ to the parameter $\theta$ of MLP and form the INR continuous function $f_{\theta}$
8:Query the corresponding value of the timestamp $C$ by $X^{\prime}=f_{\theta}(C)$ to get the reconstruct the time serial data $X^{\prime}\in\mathbb{R}^{B\times d\times T}$
9:Calculate the Anomaly Score $S\in R^{B\times T}$by measuring the point-wise difference between the input data $X$ and the reconstructed data $X^{\prime}$ as in equation (7).
10:Determine whether the point is abnormal or normal with the threshold $\delta$ as in equation (8) and get the predicted anomaly label $Y\in[0,1]^{B\times T}$
Appendix B Dataset Description
We conduct comprehensive experiments on seven multivariate datasets and one univariate dataset to validate the effectiveness of our proposed TSINR model.
Multivariate Datasets: (1) SMD (Server Machine Dataset) is a novel 5-week-long dataset collected from a large Internet company, comprising data from 28 distinct machines with 38 dimensions (Su et al., 2019). (2) PSM (Pooled Server Metrics) is sourced internally from various application server nodes at eBay with 25 dimensions and consists of 13 weeks of training data and 8 weeks of testing data (Abdulaal et al., 2021). (3) SWaT (Secure Water Treatment) is collected from an actual industrial water treatment plant and involves 51 dimensions (Mathur and Tippenhauer, 2016). (4) MSL (Mars Science Laboratory) and (5) SMAP (Soil Moisture Active Passive) are collected by spacecraft monitoring systems from NASA and comprises 55 and 25 dimensions respectively (Hundman et al., 2018, 2018). (6) PTB-XL includes clinical 12-lead Electrocardiography (ECG) that is 10 seconds in length for each patient (Jiang et al., 2023; Wagner et al., 2020). (7) SKAB is collected from the sensors installed on the testbed and includes 8 dimensions (Katser and Kozitsin, 2020).
Univariate Dataset: (8) UCR is a new benchmark introduced by multi-dataset time series anomaly detection competition in 2021 (Wu and Keogh, 2021). It encompasses 250 univariate sub-datasets, spanning diverse domains. Synthetic but highly plausible anomalies are created and each time series data in UCR has only one anomaly.
Anomaly ratio refers to the true abnormal proportion of the whole dataset while anomaly proportion indicates the abnormal proportion the user decided to choose the corresponding threshold. Anomaly proportion and window size are hyper-parameters. We follow the experimental settings of previous studies (Zhou et al., 2024; Xu et al., 2021).
Appendix C Discussion on Continuity of Multivariable Datasets
Figure 8 is a kernel density estimation plot, which is a method for visualizing the distribution of observations in a dataset. It can be observed that the SWAT dataset has the most peaks, and there is a lack of transition points between the peaks, indicating that the data changes significantly, are more high-frequency, and have poor continuity. This limitation adversely impacts the fitting capability of TSINR, consequently affecting the performance of anomaly detection.
Appendix D Study on the Group Number
Group number is a hyper-parameter in our method, representing the number of groups in the group-based architecture. In other words, it determines the number of variables learned by neurons in each group layer. Table 4 demonstrates the ablation studies on the group number. When the group number is equal to 1, it means a regular function is used rather than a group-based architecture. We observe that different datasets have different optimal parameters. This is reasonable. Drawing an analogy to image data, partitioning can accelerate the fitting of INR for images (Liu et al., 2023a). Partitioning is similar to the group-based architecture discussed here. For each image, the number of partitions is not fixed and depends on the inherent characteristics of each signal. Therefore, in our approach, the optimal group number parameter for each dataset also depends on the intrinsic features of that dataset. Nevertheless, using the group-based architecture has enhanced the model’s ability for anomaly detection.
Appendix E Study on the Anomaly Proportion
Anomaly proportion $\gamma$ is a hyper-parameter which decides the anomaly threshold $\delta$. As mentioned in Section 3.5, the threshold $\delta$ is set to label a proportion $\gamma$ of the test dataset as anomalies. We show the influence of the anomaly proportion in Table 5. It can be observed that an appropriate anomaly proportion is beneficial in aiding the model’s judgment of anomalies. Among them, PSM exhibits greater robustness to anomaly proportion compared to SMD and MSL. This is consistent with previous findings (Yang et al., 2023).
Appendix F Comparison to Other Classical Anomaly Detection Methods
Besides reconstruction-based methods, we also compared our approach with other classic anomaly detection methods, such as clustering-based methods (DeepSVDD (Ruff et al., 2018) and ITAD (Shin et al., 2020)), the density-estimation models (LOF (Breunig et al., 2000), MPPCACD (Yairi et al., 2017), DAGMM (Zong et al., 2018)), the classic methods (OCSVM (Tax and Duin, 2004), iForest (Liu et al., 2008)), and the change point detection and time series segmentation methods (BOCPD (Adams and MacKay, 2007), U-Time (Perslev et al., 2019), TS-CP2 (Deldari et al., 2021)). As shown in Table 6, TSINR achieves comparable performance on SMAP dataset and outperforms other baselines on other datasets.
Appendix G Full Experiment Results on UCR Dataset
Due to the page limitation, we place the full experimental results conducted on UCR data in this section. The UCR dataset mainly involves four domains, including human medicine, meteorology, biology and industry (Wu and Keogh, 2021). We follow the settings of the previous work (Rewicki et al., 2023) to categorize the 250 sub-datasets into four groups, with each group corresponding to a specific domain. Figure 9 illustrates the average results within each category. It can be observed that our method achieves the best performance in each domain, demonstrating its strong generalizability.
Appendix H Efficiency Analysis
We measure the efficiency of the TSINR method, and show the results in Table 7. The results indicate that the TSINR is pretty efficient and lightweight.
Figure 8. The kernel density estimation plots of five multivariate datasets.
Table 4. Ablation Studies on the group number of the group-based architecture. The best results are in Bold.
Dataset SMD PSM MSL
Metric P R F1 P R F1 P R F1
Group Num = 1 82.82 75.97 79.24 99.11 87.66 93.04 83.30 82.59 82.95
Group Num = 2 82.22 76.61 79.31 99.28 88.90 93.81 83.47 82.57 83.01
Group Num = 3 82.46 77.45 79.88 99.25 89.02 93.86 83.43 82.51 82.97
Group Num = 4 82.62 77.76 80.12 98.81 89.32 93.83 83.60 82.76 83.17
Group Num = 5 82.29 77.30 79.72 99.21 89.37 94.04 83.53 82.66 83.09
Group Num = 6 83.09 80.46 81.76 99.19 88.53 93.56 83.57 82.71 83.14
Group Num = 7 82.71 77.05 79.78 99.23 88.79 93.72 83.56 82.80 83.18
Group Num = 8 82.39 77.08 79.65 99.41 88.57 93.68 83.52 82.48 83.00
Group Num = 9 83.17 77.63 80.31 99.26 89.07 93.89 83.57 85.40 84.47
Group Num = 10 82.17 76.73 79.36 99.30 88.81 93.77 83.56 82.47 83.01
Table 5. Ablation Studies on the anomaly proportion $\gamma$ which decides the threshold $\delta$. The best results are in Bold.
Dataset SMD PSM MSL
Metric P R F1 P R F1 P R F1
$\gamma$ = 0.5 83.09 80.46 81.76 99.53 88.65 93.78 94.67 61.92 74.87
$\gamma$ = 0.6 80.55 80.93 80.74 99.37 88.90 93.85 93.16 62.36 74.71
$\gamma$ = 0.7 77.45 81.46 79.40 99.33 89.00 93.89 92.32 66.77 77.49
$\gamma$ = 0.8 75.16 83.06 78.91 99.27 88.80 93.75 91.92 73.43 81.64
$\gamma$ = 0.9 73.44 83.37 78.09 98.82 89.03 93.67 90.46 75.20 82.12
$\gamma$ = 1.0 72.55 83.39 77.59 99.21 89.37 94.04 83.98 84.26 84.12
Table 6. Comparison to other classical anomaly detection methods. The precision (P), recall (R), and F1-score (F1) values are reported, all in percentage (%). The best results are in Bold and the second-best results are in underlined.
Dataset SMD PSM SWaT MSL SMAP Average
Metric P R F1 P R F1 P R F1 P R F1 P R F1 F1
LOF 56.34 39.86 46.68 57.89 90.49 70.61 72.15 65.43 68.62 47.72 85.25 61.18 58.93 56.33 57.6 60.94
OCSVM 44.34 76.72 56.19 62.75 80.89 70.67 45.39 49.22 47.23 59.78 86.87 70.82 53.85 59.07 56.34 60.25
U-Time 65.95 74.75 70.07 82.85 79.34 81.06 46.2 87.94 60.58 57.2 71.66 63.62 49.71 56.18 52.75 65.62
IForest 42.31 73.29 53.64 76.09 92.45 83.48 49.29 44.95 47.02 53.94 86.54 66.45 52.39 59.07 55.53 61.22
DAGMM 67.3 49.89 57.3 93.49 70.03 80.08 89.92 57.84 70.4 89.6 63.93 74.62 86.45 56.73 68.51 70.18
ITAD 86.22 73.71 79.48 72.8 64.02 68.13 63.13 52.08 57.08 69.44 84.09 76.07 82.42 66.89 73.85 70.92
VAR 78.35 70.26 74.08 90.71 83.82 87.13 81.59 60.29 69.34 74.68 81.42 77.9 81.38 53.88 64.83 74.66
MMPCACD 71.2 79.28 75.02 76.26 78.35 77.29 82.52 68.29 74.73 81.42 61.31 69.95 88.61 75.84 81.73 75.74
CL-MPPCA 82.36 76.07 79.09 56.02 99.93 71.8 76.78 81.5 79.07 73.71 88.54 80.44 86.13 63.16 72.88 76.66
TS-CP2 87.42 66.25 75.38 82.67 78.16 80.35 81.23 74.1 77.5 86.45 68.48 76.42 87.65 83.18 85.36 79.00
Deep-SVDD 78.54 79.67 79.1 95.41 86.49 90.73 80.42 84.45 82.39 91.92 76.63 83.58 89.93 56.02 69.04 80.97
TSINR 83.09 80.46 81.76 99.21 89.37 94.04 99.31 72.32 83.69 83.57 85.40 84.47 91.67 76.42 83.35 85.46
Table 7. The efficiency of TSINR on the data with 128 batch size.
Training Time per Batch Inference Time per Batch Learnable Parameters
0.1s 0.08s 8.3M
Figure 9. We measure the average Precision (Left), Recall (Middle), and F1 score (Right) among different fields in the UCR datasets, which demonstrates that our TSINR is robust to time series data from different fields.